40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Nov 18 19:07:07 2011, Jenne, Update, RF System, Foam house on EOM 
    Reply  Sat Nov 19 00:09:02 2011, Zach, Update, RF System, Foam house on EOM 
Message ID: 5951     Entry time: Fri Nov 18 19:07:07 2011     Reply to this: 5954
Author: Jenne 
Type: Update 
Category: RF System 
Subject: Foam house on EOM 

[Jenne, Zach, Frank]

Frank helped Zach and I cable up at PT-100 RTD, and make sure it worked with the Newport Model 6000 Laser Diode Controller.  We're using this rather than the Newport 3040 Temperature Controller because Dmass says the output of that isn't working.  So we're using just the temp control part of the Laser Diode controller.

The back of the controller has a 15-pin D-sub, with the following useful connections.  All others were left Not Connected. 

1 & 2 (same) - Pin 2 is one side of TEC output (we have it connected to one side of a resistive heater)

3 & 4 (same) - Pin 4 is the other side of the TEC output (connected to the other side of the resistive heater)

7 - connected to one side of PT-100 temp sensor

8 - connected to other side of PT-100 temp sensor

I used aluminum tape to attach the sensor and heater to the 40m's EOM, and we plugged in the controller.  It seems to be kind of working.  Zach figured out the GPIB output stuff, so we can talk to it remotely.

 

ELOG V3.1.3-